The value of serum sodium in healthy adults approximately follows a normal distribution with a mean of 141meq/l and standard deviation of 3meq/l


(a) What is the probability that a normal healthy adult will have a serum sodium value above 147meq/l ( 5 marks)


(b) What is the probability that a normal healthy adult will have a serum sodium value below 130 meq/l ( 5 marks)


(c) What is the probability that a normal healthy adult will have a serum sodium value between 132 and 150meq/l (5 marks)

1
Expert's answer
2022-04-18T00:11:42-0400

We have a normal distribution,Β "\mu=141, \sigma=3."

Let's convert it to the standard normal distribution,

"z=\cfrac{x-\mu}{\sigma}."


"(a) \ z=\cfrac{147-141}{3}=2;\\
P(X>147)=P(Z>2)=\\
=1-P(Z<2)=\\
=1-0.9772=0.0228 \text{ (from z-table).}"


"(b) \ z=\cfrac{130-141}{3}=-3.67;\\
P(X<130)=P(Z<-3.67)=\\
=0.0001 \text{ (from z-table).}"


"(c) \ z_1=\cfrac{132-141}{3}=-3;\\
z_2=\cfrac{150-141}{3}=3;\\
P(132<X<150)=P(-3<Z<3)=\\
=P(Z<3)-P(Z<-3)=\\
=0.9987-0.0013=0.9974 \text{ (from z-table).}"